OpenStudy (anonymous):

The following table shows the values of f(x) and g(x) for different values of x: x 1 2 3 4 5 f(x) 3 9 27 81 243 g(x) 4 4 4 4 4 Which of the following best describes the function f(x) + g(x)? 3x + 4 3x + 4 3x + 4x 4(3x)

OpenStudy (solomonzelman):

can you please write the tables normally ?

OpenStudy (anonymous):

yeah sorry x 1 2 3 4 5 f(x) 3 9 27 81 243 g(x) 4 4 4 4 4

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (kohai):

The first two options look exactly the same

OpenStudy (kohai):

I'm sure one is supposed to be different, please recheck them

OpenStudy (kohai):

All the same, I can help you with this question if you still need it

OpenStudy (kohai):

The answer is, 3x+4. I'll explain each section to you. 3x, It asks for f(x)+g(x). The f(x) begins with 3, and if a number is put in, it'll increase as you see on the table. +4, Each value on the g(x) is 4, so there won't be any x afterwards because it's not increasing.

OpenStudy (anonymous):

oh sorry i have to right them on the equation thing \[3^x +4\] then 3^X^+4 (couldent figure out how to do it on the equation thing) \[3^x+4x\] \[4(3^x)\] those are the choices @kohai

OpenStudy (kohai):

It'll be the first one, in that case. Take a look, 3^1=3 3^2=3*3=9 3^3=3*3*3=27 That follows the table correctly. As far as the g(x) goes, it's 4 for each value so there won't be an x. Also, it asks for f(x)+g(x), so you can scratch a few answer choices off that way

OpenStudy (anonymous):

lol thanks do you think you can help me with a few more im not going to give you like 20 more like other people do just 4

OpenStudy (kohai):

Close this question and then post a new one, I'll take a look at it. If I can help you I will, if not it'll give other people an opportunity to look at it

OpenStudy (anonymous):

kk